# Volume and topology of bounded and closed hyperbolic $3$-manifolds, II

### v1

@@ 1. Introduction
[…]
@@
[…]
@@
The proofs of Theorems 10.1 and 11.3 and Proposition 9.4 involve combining a rich variety of topological and geometric techniques. One key ingredient in the proof of Proposition 9.4 is the following result, which pertains to manifolds with boundary and features a different measure of topological complexity, the Heegaard genus, which in this paper will be denoted $\mathrm{Hg}(N)$. For a compact, connected, orientable $3$-manifold $N$, with or without boundary, ${\rm Hg}(N)$ is the smallest genus of a closed surface in $N$ that divides it into two compression bodies.
@@
This directly strengthens Theorem 1.1 of [14], which gives a lower bound of $6.89$ on the volume of $N$ with the same topological hypothesis. Like the results above for closed manifolds, it is still likely not sharp: in a census compiled by Frigerio-Martelli-Petronio [17], all examples with volume at most $2V_{\rm oct}$ have Heegaard genus $3$. This census collected all manifolds with totally geodesic boundary that decompose into at most four truncated tetrahedra. After Kojima–Miyamoto’s minimal-volume examples (of volume $6.45$), the census of [17] contains six manifolds with volume $7.10$, and the next-smallest have volume $7.33...~{}>2V_{\rm oct}$.
@@
Section 2 reviews the methods introduced in [27] and further developed in [14] for describing how the lengths of return paths of a manifold $N$ satisfying the hypotheses of Theorem 7.2 are controlled by the geometry of $\partial N$, moderated by $(i,j,k)$-hexagons. In places, we incrementally improve these methods. In particular, Proposition 2.2.2 simplifies the lower bound on the second-shortest return path length $\ell_{2}$ as a function of $\ell_{1}$ given in [14, Lemma 2.9] (which itself built on [27, §4]). And Proposition 2.3.1 gives a stronger lower bound than [14, Prop. 3.9] on $\cosh\ell_{1}$ in the absence of a $(1,1,1)$-hexagon: $1.23$ here versus $1.215$ there. (The minimum possible value for $\cosh\ell_{1}$ is $1.183$, proved in [27].)
@@
In the paper’s third part, beginning with Section 8, we shift our focus to closed manifolds. Section 8 gives background necessary to prove our results in this setting. The proof of Proposition 9.4 begins by using topological results due to Culler and Shalen [11], about desingularization of $\pi_{1}$-injective singular surfaces in $3$-manifolds, to show that if $M$ satisfies the homological hypothesis of the proposition and $\pi_{1}(M)$ is not $k$-free, then $M$ contains a closed incompressible surface $S_{0}$ of some genus $g$ with $12\leq g\leq k-$ 1. The homological hypothesis implies that the Heegaard genus of $M$ is strictly greater than $12g+$ 1.
@@
The existence of an incompressible surface $S_{0}$ of genus $g$, together with the strict lower bound $12g+$ 1 for the Heegaard genus of $M$, is used—via a result proved in [18] using topological ideas developed by Culler, DeBlois and Shalen in [8]—to produce a closed incompressible surface $S\subset M$ such that either (1) the Euler characteristic $\chi(\mathop{\rm kish}(M\setminus\setminus S))$ is at most $-2$, or (2) $S$ separates $M$, and $M\setminus\setminus S$ has a component which is acylindrical, i.e. contains no essential annulus. If (1) holds, the geometric methods developed by Agol, Storm and Thurston in [4] (via Theorem 6.2.8 of this paper) give a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}M$.
If (2) holds but (1) does not, and if we fix an acylindrical component $A$ of $M\setminus\setminus S$, then $A$ is diffeomorphic to a hyperbolic $3$-manifold $N$ with totally geodesic boundary, and the methods of [4] show that $\mathop{\rm vol}N$ is a lower bound for $\mathop{\rm vol}M$. Furthermore, in this case a Mayer-Vietoris calculation shows that $\dim H_{1}(N,{\mathbb{F}}_{2})$, and hence the Heegaard genus of $N$, is at least $5$. If $S$ has genus $2$, Theorem 7.2 now gives a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}N$. If $S$ has genus greater than $2$, the geometric results established by Miyamoto in [30] give a stronger lower bound.
To prove Theorem 10.1 one must show that if that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 5$ then $\mathop{\rm vol}M>V_{\rm oct}/2$. If $\pi_{1}(M)$ is $3$-free, then results proved by Anderson, Canary, Culler and Shalen in [5], and improved in [3] by using the celebrated tameness theorem proved by Agol in of [1] and by Calegari-Gabai in [7], give a lower bound for $\mathop{\rm vol}M$ of $3.08$, which is considerably bigger than $V_{\rm oct}/2$. If $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 5$ but $\pi_{1}(M)$ is not $3$-free, a novel but simple application of one of the topological results established by Shalen and Wagreich in [34] provides a $({\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}})$-covering space $\widetilde{M}$ of $M$ such that $\pi_{1}(\widetilde{M})$ is not $3$-free, and $\dim H_{1}(\widetilde{M};{\mathbb{F}}_{2})\geq 7$. One can then use Proposition 9.4 to show that $\mathop{\rm vol}\widetilde{M}>2V_{\rm oct},$ which implies $\mathop{\rm vol}M>V_{\rm oct}/2$.
@@
If $M$ satisfies the hypothesis of Assertion (1) of (2) of Theorem 11.3 and $\pi_{1}(M)$ is not $4$-free, then one uses the homological hypotheses to find a two-sheeted covering $\widetilde{M}$ of $M$ such that $\pi_{1}(\widetilde{M})$ is not $4$-free, and $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 8$. Proposition 9.4 then provides a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}\widetilde{M}$, and hence a strict lower bound of $V_{\rm oct}$ for $\mathop{\rm vol}M$. (This is similar to the argument used to prove [14, Theorem 1.2] in the non-free case, but we obtain a stronger lower bound in this context thanks to Theorem 7.2. In our proof of Theorem 11.3, in addition to improving the estimates given by Theorem 1.2 of [14], we have taken the opportunity to provide more detail than was given in the proof of the latter result, and to correct a citation that appeared in that proof.)
@@ 2. Existing foundations
This section reviews an approach developed by Kojima–Miyamoto [27], and further exploited in our earlier paper [14], to controlling the geometry of a hyperbolic $3$-manifold $N$ with totally geodesic boundary via the geometry of $\partial N$. We focus on the orthospectrum of $N$, the sequence of lengths of properly immersed arcs in $N$ that meet $\partial N$ perpendicularly.
@@ 2.1. Ortholengths vs arclengths
For a finite-volume hyperbolic $3$-manifold $N$ with compact totally geodesic boundary and three components $\Pi_{1}$, $\Pi_{2}$, and $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the right-angled hexagon supplied by Lemma 2.1.4. We call the edges of $C$ which are short cuts internal, and the edges in $\partial\widetilde{N}$ external. If the internal edges lift $\lambda_{i}$, $\lambda_{j}$, and $\lambda_{k}$, we call $C$ an $(i,j,k)$ hexagon.
@@
For a finite-volume hyperbolic $3$-manifold $N$ with compact totally geodesic boundary and any fixed $i,j\in\mathbb{N}$, if $k_{1}\leq k_{2}\leq k_{3}\ldots$ is the set of $k\in\mathbb{N}$ such that there exists an $(i,j,k)$ hexagon in $\widetilde{N}$ then $d_{ij}=X_{ij}^{k_{1}}$, and $d_{ij}^{(n)}=X_{ij}^{k_{n}}$ for $n>1$. In particular, $d_{ij}^{k}\geq X_{ij}^{k}$ for all $i,j,k\in\mathbb{N}$.
@@ 2.2. Kojima–Miyamoto’s lower bound
The basic idea here is that for a hyperbolic $3$-manifold $N$ with compact totally geodesic boundary, the topology of $\partial N$ determines its area by the Gauss–Bonnet theorem, and this bounds the areas of disks in a packing of $\partial N$. The radii of such a packing are determined by the orthospectrum of $N$.
@@
and let $S=X_{12}^{1}-R$. For a finite-volume hyperbolic $3$-manifold $N$ with compact, connected totally geodesic boundary such that $S>0$ there are four disks embedded in $\partial N$ without overlapping: two of radius $R$, centered at feet of $\lambda_{1}$, and two of radius $S$, centered at feet of $\lambda_{2}$.
@@
For a finite-volume hyperbolic $3$-manifold $N$ with compact, connected totally geodesic boundary of genus $2$, $\ell_{2}\geq\max\{\ell_{1},E,M\}$.
@@ 2.3. A bound for $\ell_{1}$ in the absence of a $(1,1,1)$-hexagon
Proposition 3.9 of [14] bounds the first ortholength $\ell_{1}$ of a finite-volume hyperbolic $3$-manifold $N$ with compact totally geodesic boundary below by $\cosh\ell_{1}\geq 1.215$, assuming $\widetilde{N}$ has no $(1,1,1)$-hexagon. This is significantly better than the sharp universal lower bound of $\frac{3+\sqrt{3}}{4}\simeq 1.183$ for $\cosh\ell_{1}$ proved in [27]. (For comparison, all manifolds with volume less than $7.63$ in the census of Petronio et. al. have $\cosh\ell_{1}<1.213$.)
@@
For the absolute lower bound on $\cosh\ell_{1}$, we note that the upper bound we have just proved for $\ell_{2}$ is an increasing function of $\ell_{1}$, whereas the lower bound $E$ of Proposition 2.2.2 is decreasing. Direct computation shows that $E$ takes the value $1.200$ and the upper bound the value $1.194$ when $\cosh\ell_{1}=1.23$ (and their values coincide when $\cosh\ell_{1}\approx 1.2304$). Thus if there is no $(1,1,1)$-hexagon and $\cosh\ell_{1}\leq 1.23$, the lower bound on $\ell_{2}$ exceeds the upper bound, a contradiction.∎
@@ 3. More Better Muffins
A “muffin”, so named by Kojima and Miyamoto in [27] (see the discussion below Proposition 3.1 of that paper), is a member of a certain class of hyperbolic solids of rotation, shown in [27] to embed in a compact hyperbolic three-manifold with totally geodesic boundary and used in the main volume bound of [27]. We used the same class of muffins from [27] in [14], denoting them as “$\mathrm{Muf}_{\ell_{1}}$” in Definition 3.1 there. As this notation suggests, members of this class are determined up to isometry by a single parameter, which in their application to volume bounds is the length $\ell_{1}$ of the shortest return path. Here we will continue to use the notation of [14] for this class of muffins (see Definitions and Remarks 3.1.6 below).
This section introduces a more general class of muffins, still hyperbolic solids of rotation, but now depending on two parameters $\ell$ and $R$ which are side lengths of certain reflectively symmetric hyperbolic pentagons. In Section 4 we will use muffins together with collars to give lower bounds on volume for a hyperbolic $3$-manifold $N$ with totally geodesic boundary satisfying certain bounds on the lengths of $\ell_{1}$ and $\ell_{2}$. Here we formally define these objects and establish technical results — for muffins in Section 3.1 and collars in Section 3.2 — that will allow us to show that they are disjointly embedded in $N$.
@@ 3.1. Muffins
Given $\ell,R>0$, let $Q$ be a hyperbolic Lambert quadrilateral—one with three right angles—such that its sides having right angles at both endpoints are of lengths $\ell/2$ and $R$. Doubling $Q$ across its edge $\omega$ opposite the one $\rho$ with length $R$ yields a reflectively symmetric hyperbolic pentagon $P$ with four right angles. The base $\lambda$ of $P$—the side opposite the non-right vertex—has length $\ell$, and each of the two sides intersecting it have length $R$. A muffin is the solid in $\mathbb{H}^{3}$ obtained by rotating $P$ around $\lambda$. We illustrate this construction in Figure 3.1 and formalize its definition below.
@@
For a given hyperbolic manifold $N$ with totally geodesic boundary, we will embed a muffin in $N$—for carefully chosen $\ell$ and $R$—by checking that a copy of $M(\ell,R)$ carefully placed in the universal cover $\widetilde{N}$ does not intersect its translates under the action of $\pi_{1}(N)$ by covering transformations. Here as in the prior works [27] and [14], “carefully placed” will always mean centered at a lift $\tilde{\lambda}$ of a return path $\lambda$, with height $\ell$ equal to the length of $\lambda$ so that each of its caps lies in a component of $\partial\widetilde{N}$. We use the following result to ensure that $M(\ell,R)$, with choices carefully made, lies in $\widetilde{N}$.
@@
For $\tilde{\lambda}_{k}$ as above, let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing its endpoints. For a third component $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the planar right-angled hexagon supplied by Lemma 2.1.4, containing $\tilde{\lambda}_{k}$ and intersecting each $\Pi_{i}$ perpendicularly in a side, for $i=1$, $2$, $3$. Then $C\cap\Pi_{3}$ is the side of $C$ opposite $\tilde{\lambda}_{k}$.
@@
The geometric motivation for the choice of cap radius for $\mathrm{Muf}_{\ell_{1}}$ is captured in Figure 3.1 of [27]. The left side of that Figure pictures a $(1,1,1)$-hexagon in our notation, with “$\ell$” there equal to $\ell_{1}$ and “$R$” and “$A$” given by (2.2.1.1) and (3.1.6.1). The Fact below records a feature of that picture.
@@
For a hyperbolic $3$-manifold $N$ and a lift $\tilde{\lambda}_{k}$ of the $k$th-shortest return path of $N$ to the universal cover $\widetilde{N}$, a copy of $M(\ell_{k},R)$ centered at $\tilde{\lambda}_{k}$ is contained in $\widetilde{N}$ if its waist radius is less than the quantity $U_{k}$ defined in Lemma 3.1.5. If so, it intersects $\partial\widetilde{N}$ precisely in its caps.
@@
Let =$\mathcal{B}=\{\Pi_{1},\Pi_{2},\Pi_{1}^{\prime},\Pi_{2}^{\prime}\}$, and for $i=1,2$, let $H_{i}$ (respectively, $H_{i}^{\prime}$) be the half-space bounded by $\Pi_{i}$ (resp. $\Pi_{i}^{\prime}$) that contains $\widetilde{N}$ and hence the other $\Pi_{j}$ and $\Pi_{j}^{\prime}$. Recall from Lemma 2.1.3 that each three-element subset of $\mathcal{B}$ determines a geodesic plane that is perpendicular to all three of its members. It follows from Lemma 3.1 of [13] that either all four planes above have a common perpendicular plane $\Pi$, or that the common perpendicular to each three-plane subcollection bounds a single half-space that contains all three lifted return paths joining its members to the fourth plane. The truncated tetrahedron $\Delta$ determined by $\mathcal{B}$ is described in [13, Definition 3.2]. In the latter case, it is the intersection of the $H_{i}$ and $H_{i}^{\prime}$ with the four half-spaces determined in this way by the three-element subcollections of $\mathcal{B}$. In the former, we take $\Delta$ to be the intersection of the common perpendicular plane $\Pi$ with the $H_{i}$ and $H_{i}^{\prime}$, and say it is degenerate.
@@
Since $23x-1<4x-$ 2 for $x>1$ and $y\mapsto\tanh y$ is increasing, it follows that $D/2>W$, hence that $M$ does not intersect $M^{\prime}$. ∎
@@
The right-hand side above vanishes at $x=1$, but we are interested only in $x>1$. Setting the right-hand side above equal to 00 and solving for $y$ yields
@@
[…]
@@
We claim first that $W_{S}^{0}$ is an upper bound for the waist radius $W_{2}$ of $M(\ell_{2},S)$, taken as a function of $x$ and $y$, on the rectangle $[a,b]\times[c,d]$. By the formula (3.1.1.1), $W_{2}$ satisfies $\tanh W_{2}=\cosh(\ell_{2}/2)\tanh S$. Lemma 3.1.10 implies that as a function of $x$ and $y$, $S$ increases with $x$ and decreases with $y$ at all points of this rectangle. The values of $S$ are thus bounded above by $S(b,c)$ on $[a,b]\times[c,d]$, and the claim thus follows from the “half-angle identity” for hyperbolic cosine.

### v2

@@ 1. Introduction
Among these results, only Theorem 1.1 of [2] is known to be sharp, and only for $p=5$: we have $\dim H_{1}(M;{\mathbb{F}}_{5})=2$ when $M$ is the Weeks manifold. Among the closed, orientable hyperbolic $3$-manifolds $M$ of volume at most $3.69$ known from a census compiled by Hodgson–Weeks (see [23]) and shipped with SnapPy [9], the largest value of $\dim H_{1}(M;{\mathbb{F}}_{2})$ is $3$; it is achieved by the manifold denoted by ${\rm m135}(-1,3)$ in the census,
@@
This result improves on Theorem 1.2 of [11] in the range in which it applies, but it is probably not sharp. Among known closed, orientable hyperbolic $3$-manifolds $M$ of volume at most $V_{\rm oct}/2$, the largest value of $\dim H_{1}(M;{\mathbb{F}}_{2})$ is $2$; it is achieved by the manifold denoted by ${\rm m}009(5,1)$ in the census cited above, which is arithmetic and has volume exactly $V_{\rm oct}/2$.
@@
The proofs of Theorems 10.1 and 11.3 and Proposition 9.4 involve combining a rich variety of topological and geometric techniques. One key ingredient in the proof of Proposition 9.4 is the following result, which pertains to manifolds with boundary and features a different measure of topological complexity, the Heegaard genus, which in this paper will be denoted $\mathrm{Hg}(N)$. For a compact, connected, orientable $3$-manifold NN, with or without boundary, ${\rm Hg}(N)$ is the smallest genus of a closed surface in $N$ that divides it into two compression bodies.
@@
This directly strengthens Theorem 1.1 of [14], which gives a lower bound of $6.89$ on the volume of $N$ with the same topological hypothesis. Like the results above for closed manifolds, it is still likely not sharp: in a census compiled by Frigerio-Martelli-Petronio [17], all examples with volume at most $2V_{\rm oct}$ have Heegaard genus $3$. This census collected all manifolds with totally geodesic boundary that decompose into at most four truncated tetrahedra. After Kojima–Miyamoto’s minimal-volume examples (of volume $6.45$...), the census of [17] contains six manifolds with volume $7.10$..., and the next-smallest have volume $7.33...~>2V_{\rm oct}$.
@@
Section 2 reviews the methods introduced in [27] and further developed in [14] for describing how the lengths of return paths of a manifold $N$ satisfying the hypotheses of Theorem 7.2 are controlled by the geometry of $\partial N$, moderated by $(i,j,k)$-hexagons. In places, we incrementally improve these methods. In particular, Proposition 2.2.2 simplifies the lower bound on the second-shortest return path length $\ell_{2}$ as a function of $\ell_{1}$ given in [14, Lemma 2.9] (which itself built on [27, §4]). And Proposition 2.3.1 gives a stronger lower bound than [14, Prop. 3.9] on $\cosh\ell_{1}$ in the absence of a $(1,1,1)$-hexagon: $1.23$ here versus $1.215$ there. (The minimum possible value for $\cosh\ell_{1}$ is $1.183..$, proved in [27].)
@@
In the paper’s third part, beginning with Section 8, we shift our focus to closed manifolds. Section 8 gives background necessary to prove our results in this setting. The proof of Proposition 9.4 begins by using topological results due to Culler and Shalen [11], about desingularization of $\pi_{1}$-injective singular surfaces in $3$-manifolds, to show that if $M$ satisfies the homological hypothesis of the proposition and $\pi_{1}(M)$ is not $k$-free, then $M$ contains a closed incompressible surface $S_{0}$ of some genus $g$ with $2\leq g\leq k-1$. The homological hypothesis implies that the Heegaard genus of $M$ is strictly greater than $2g+1$.
@@
The existence of an incompressible surface $S_{0}$ of genus $g$, together with the strict lower bound $2g+1$ for the Heegaard genus of $M$, is used—via a result proved in [18] using topological ideas developed by Culler, DeBlois and Shalen in [8]—to produce a closed incompressible surface $S\subset M$ such that either (1) the Euler characteristic $\chi(\mathop{\rm kish}(M\setminus\setminus S))$ is at most 2-2, or (2) $S$ separates $M$, and $M\setminus\setminus S$ has a component which is acylindrical, i.e. contains no essential annulus. If (1) holds, the geometric methods developed by Agol, Storm and Thurston in [4] (via Theorem 6.2.8 of this paper) give a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}M$.
If (2) holds but (1) does not, and if we fix an acylindrical component $A$ of $M\setminus\setminus S$, then $A$ is diffeomorphic to a hyperbolic $3$-manifold NN with totally geodesic boundary, and the methods of [4] show that $\mathop{\rm vol}N$ is a lower bound for $\mathop{\rm vol}M$. Furthermore, in this case a Mayer-Vietoris calculation shows that $\dim H_{1}(N,{\mathbb{F}}_{2})$, and hence the Heegaard genus of $N$, is at least $5$. If $S$ has genus $2$, Theorem 7.2 now gives a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}N$. If $S$ has genus greater than $2$, the geometric results established by Miyamoto in [30] give a stronger lower bound.
[…]
@@
We are grateful to Nathan Dunfield for explaining the example ${\rm m}009(5,1)$ that was referred to above, and to Joel Hass for explaining material related to Proposition 6.2.7 to us.
@@ 2. Existing foundations
This section reviews an approach developed by Kojima–Miyamoto [27], and further exploited in our earlier paper [14], to controlling the geometry of a hyperbolic $3$-manifold NN with totally geodesic boundary via the geometry of $\partial N$. We focus on the orthospectrum of $N$, the sequence of lengths of properly immersed arcs in $N$ that meet $\partial N$ perpendicularly.
@@ 2.1. Ortholengths vs arclengths
For a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary and three components $\Pi_{1}$, $\Pi_{2}$, and $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the right-angled hexagon supplied by Lemma 2.1.4. We call the edges of $C$ which are short cuts internal, and the edges in $\partial\widetilde{N}$ external. If the internal edges lift $\lambda_{i}$, $\lambda_{j}$, and $\lambda_{k}$, we call $C$ an $(i,j,k)$ hexagon.
@@
For a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary and any fixed $i,j\in\mathbb{N}$, if $k_{1}\leq k_{2}\leq k_{3}\ldots$ is the set of $k\in\mathbb{N}$ such that there exists an $(i,j,k)$ hexagon in $\widetilde{N}$ then $d_{ij}=X_{ij}^{k_{1}}$, and $d_{ij}^{(n)}=X_{ij}^{k_{n}}$ for $n>1$. In particular, $d_{ij}^{k}\geq X_{ij}^{k}$ for all $i,j,k\in\mathbb{N}$.
@@ 2.2. Kojima–Miyamoto’s lower bound
The basic idea here is that for a hyperbolic $3$-manifold NN with compact totally geodesic boundary, the topology of $\partial N$ determines its area by the Gauss–Bonnet theorem, and this bounds the areas of disks in a packing of $\partial N$. The radii of such a packing are determined by the orthospectrum of $N$.
@@
and let $S=X_{12}^{1}-R$. For a finite-volume hyperbolic $3$-manifold NN with compact, connected totally geodesic boundary such that $S>0$ there are four disks embedded in $\partial N$ without overlapping: two of radius $R$, centered at feet of $\lambda_{1}$, and two of radius $S$, centered at feet of $\lambda_{2}$.
@@
For a finite-volume hyperbolic $3$-manifold NN with compact, connected totally geodesic boundary of genus $2$, $\ell_{2}\geq\max\{\ell_{1},E,M\}$.
@@ 2.3. A bound for $\ell_{1}$ in the absence of a $(1,1,1)$-hexagon
Proposition 3.9 of [14] bounds the first ortholength $\ell_{1}$ of a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary below by $\cosh\ell_{1}\geq 1.215$, assuming $\widetilde{N}$ has no $(1,1,1)$-hexagon. This is significantly better than the sharp universal lower bound of $\frac{3+\sqrt{3}}{4}\simeq 1.183$ for $\cosh\ell_{1}$ proved in [27]. (For comparison, all manifolds with volume less than $7.63$ in the census of Petronio et. al. have $\cosh\ell_{1}<1.213$.)
@@
For the absolute lower bound on $\cosh\ell_{1}$, we note that the upper bound we have just proved for $\ell_{2}$ is an increasing function of $\ell_{1}$, whereas the lower bound $E$ of Proposition 2.2.2 is decreasing. Direct computation shows that $E$ takes the value $1.200$... and the upper bound the value $1.194$... when $\cosh\ell_{1}=1.23$ (and their values coincide when $\cosh\ell_{1}\approx 1.2304$). Thus if there is no $(1,1,1)$-hexagon and $\cosh\ell_{1}\leq 1.23$, the lower bound on $\ell_{2}$ exceeds the upper bound, a contradiction.∎
@@ 3. More Better Muffins
A “muffin”, so named by Kojima and Miyamoto in [27] (see the discussion below Proposition 3.1 of that paper), is a member of a certain class of hyperbolic solids of rotation, shown in [27] to embed in a compact hyperbolic three-manifold with totally geodesic boundary and used in the main volume bound of [27]. We used the same class of muffins from [27] in [14], denoting them as $\mathrm{Muf}_{\ell_{1}}$” in Definition 3.1 there. As this notation suggests, members of this class are determined up to isometry by a single parameter, which in their application to volume bounds is the length $\ell_{1}$ of the shortest return path. Here we will continue to use the notation of [14] for this class of muffins (see Definitions and Remarks 3.1.6 below).
This section introduces a more general class of muffins, still hyperbolic solids of rotation, but now depending on two parameters $\ell$ and $R$ which are side lengths of certain reflectively symmetric hyperbolic pentagons. In Section 4 we will use muffins together with collars to give lower bounds on volume for a hyperbolic $3$-manifold NN with totally geodesic boundary satisfying certain bounds on the lengths of $\ell_{1}$ and $\ell_{2}$. Here we formally define these objects and establish technical results — for muffins in Section 3.1 and collars in Section 3.2 — that will allow us to show that they are disjointly embedded in $N$.
@@ 3.1. Muffins
Given $\ell,R>0$, let $Q$ be a hyperbolic Lambert quadrilateral—one with three right angles—such that its sides having right angles at both endpoints are of lengths $\ell/2$ and $R$. Doubling $Q$ across its edge $\omega$ opposite the one $\rho$ with length $R$ yields a reflectively symmetric hyperbolic pentagon $P$ with four right angles. The base $\lambda$ of PP—the side opposite the non-right vertex—has length $\ell$, and each of the two sides intersecting it have length $R$. A muffin is the solid in $\mathbb{H}^{3}$ obtained by rotating $P$ around $\lambda$. We illustrate this construction in Figure 3.1 and formalize its definition below.
@@
For a given hyperbolic manifold $N$ with totally geodesic boundary, we will embed a muffin in NN—for carefully chosen $\ell$ and RR—by checking that a copy of $M(\ell,R)$ carefully placed in the universal cover $\widetilde{N}$ does not intersect its translates under the action of $\pi_{1}(N)$ by covering transformations. Here as in the prior works [27] and [14], “carefully placed” will always mean centered at a lift $\tilde{\lambda}$ of a return path $\lambda$, with height $\ell$ equal to the length of $\lambda$ so that each of its caps lies in a component of $\partial\widetilde{N}$. We use the following result to ensure that $M(\ell,R)$, with choices carefully made, lies in $\widetilde{N}$.
@@
For $\tilde{\lambda}_{k}$ as above, let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing its endpoints. For a third component $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the planar right-angled hexagon supplied by Lemma 2.1.4, containing $\tilde{\lambda}_{k}$ and intersecting each $\Pi_{i}$ perpendicularly in a side, for $i=1$, 22, 33. Then $C\cap\Pi_{3}$ is the side of $C$ opposite $\tilde{\lambda}_{k}$.
@@
The geometric motivation for the choice of cap radius for $\mathrm{Muf}_{\ell_{1}}$ is captured in Figure 3.1 of [27]. The left side of that Figure pictures a $(1,1,1)$-hexagon in our notation, with $\ell$” there equal to $\ell_{1}$ and “RR” and “AA” given by (2.2.1.1) and (3.1.6.1). The Fact below records a feature of that picture.
@@
For a hyperbolic $3$-manifold NN and a lift $\tilde{\lambda}_{k}$ of the $k$th-shortest return path of $N$ to the universal cover $\widetilde{N}$, a copy of $M(\ell_{k},R)$ centered at $\tilde{\lambda}_{k}$ is contained in $\widetilde{N}$ if its waist radius is less than the quantity $U_{k}$ defined in Lemma 3.1.5. If so, it intersects $\partial\widetilde{N}$ precisely in its caps.
@@
Let $\mathcal{B}=\{\Pi_{1},\Pi_{2},\Pi_{1}^{\prime},\Pi_{2}^{\prime}\}$, and for $i=1,2$, let $H_{i}$ (respectively, $H_{i}^{\prime}$) be the half-space bounded by $\Pi_{i}$ (resp. $\Pi_{i}^{\prime}$) that contains $\widetilde{N}$ and hence the other $\Pi_{j}$ and $\Pi_{j}^{\prime}$. Recall from Lemma 2.1.3 that each three-element subset of $\mathcal{B}$ determines a geodesic plane that is perpendicular to all three of its members. It follows from Lemma 3.1 of [13] that either all four planes above have a common perpendicular plane $\Pi$, or that the common perpendicular to each three-plane subcollection bounds a single half-space that contains all three lifted return paths joining its members to the fourth plane. The truncated tetrahedron $\Delta$ determined by $\mathcal{B}$ is described in [13, Definition 3.2]. In the latter case, it is the intersection of the $H_{i}$ and $H_{i}^{\prime}$ with the four half-spaces determined in this way by the three-element subcollections of $\mathcal{B}$. In the former, we take $\Delta$ to be the intersection of the common perpendicular plane $\Pi$ with the $H_{i}$ and $H_{i}^{\prime}$, and say it is degenerate.
@@
Since 3x1<4x23x-1<4x-2 for $x>1$ and $y\mapsto\tanh y$ is increasing, it follows that $D/2>W$, hence that $M$ does not intersect $M^{\prime}$. ∎
@@
The right-hand side above vanishes at $x=1$, but we are interested only in $x>1$. Setting the right-hand side above equal to 0 and solving for $y$ yields
@@
(3.1.11.1) $T_{12}^{0}=\cosh^{-1}\left(\frac{2b}{\sqrt{(b-1)(d-1)}}\right),$ $\mbox{and}\quad T_{22}^{0}=\cosh^{-1}\left(\frac{2a}{d-1}\right),$
@@
We claim first that $W_{S}^{0}$ is an upper bound for the waist radius $W_{2}$ of $M(\ell_{2},S)$, taken as a function of $x$ and $y$, on the rectangle $a,b]\times[c,d]$. By the formula (3.1.1.1), $W_{2}$ satisfies $\tanh W_{2}=\cosh(\ell_{2}/2)\tanh S$. Lemma 3.1.10 implies that as a function of $x$ and $y$, $S$ increases with $x$ and decreases with $y$ at all points of this rectangle. The values of $S$ are thus bounded above by $S(b,c)$ on $a,b]\times[c,d]$, and the claim thus follows from the “half-angle identity” for hyperbolic cosine.